2. A sample of 40 sales receipts from a grocery store has sample mean of 137 and population sd= 30.2. Use these values to test whether or not the mean sales the grocery store is less than 150. Use a 95% confidence level

1
Expert's answer
2022-06-06T03:22:30-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge150"

"H_1:\mu<150"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=39" and the critical value for a left-tailed test is "t_c =-1.684875."

The rejection region for this left-tailed test is "R = \{t:t<-1.684875\}."

The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{137-150}{30.2/\sqrt{40}}=-2.7225"


Since it is observed that "t=-2.7225<-1.684875=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed, "df=39" degrees of freedom, "t=-2.7225" is "p=0.004819," and since "p=0.004819<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 150, at the "\alpha = 0.05" significance level.